OpenStudy (anonymous):

I need help with factoring a^2+81, I come up with a^2+9^2=(a+3)(a+3) but it is wrong??

OpenStudy (anonymous):

You cannot factorise this. there is nothing like the sum of two squares!!!!!!!!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

general rule when the discriminant of a quadratic is negative u cant factorize it. here u have\[\Delta=0^2-4\times81<0\]

OpenStudy (precal):

your question is prime because of the + sign

OpenStudy (shane_b):

@precal: \[\text{In your first image it should be }a^2-b^2=(a+b)(a-b)\]

OpenStudy (precal):